### v2

@@ 4.2 Proof of Theorem 5
Consider the case where $f,g$ are weighted contracts and suppose towards a contradiction that $p\in E(f)\cap E(g)$. Let $i\in[n]$ denote the agent with the smallest weight ratio $\frac{\lambda^{g}_{i}}{\lambda_{i}^{f}}$, where $\lambda^{f},\lambda^{g}$ represent the weights that define the contracts $f,g\in\mathcal{F}_{W}$. It follows from the definition of weighted contracts that for any $S\subset[n]$ such that $i\in S$, agent $i$’s reward is weakly lower under $g$ than $f$. As a result, it must be that $r_{i}(g,p_{-i})<r_{i}(f,p_{-i})$ and so it can’t be that $b_{i}(g,p_{-i})=b_{i}(f,p_{-i})$. It follows that two different weighted contracts cannot have the same equilibrium.
Moreover, for any $j\in T^{\prime}\setminus T$,
Thus,
However, since $p$ is an equilibrium for $f$ and $g$, both of the outside expressions must be equal to $\sum_{k\in T}p_{k}\cdot c_{k}^{\prime}(p_{k})$, contradiction. ∎
To conclude this section, in the following Corollary, we characterize the set of maximal equilibria in terms of the model parameters.
Corollary 4.
$p\in\mathcal{P}$ if and only if $p$ satisfies (5) and $z(p)=1$.
Thus, the principal can potentially try to optimize its objective $V(p)$ over the set of solutions to the equation $z(p)=1$, and then check if the solution satisfies (5).
5 Discussion
In this section, we briefly discuss how our results extend to environments where the small budget assumption may not hold, and also discuss the desirability of Luce contracts for implementation over other commonly-studied contracts like piece-rate contracts or bonus-pool contracts in standard multi-agent contract design environments.
Larger budgets
In this subsection, we discuss how our results extend to environments where the small budget assumption may not hold, so that $B>\min\{c_{1}^{\prime}(1),\dots,c_{n}^{\prime}(1)\}$. Under the small budget assumption, we show that equilibria are maximal if and only if they are implementable exclusively by SGE contracts (Theorem 3.2). While the result does not exactly extend to environments where the budget is large, it extends as long as we restrict attention to equilibria that are interior. Indeed, the small budget assumption ensures that all $p\in\mathcal{E}$ are such that $p_{i}<1$ for all $i\in[n]$, which is necessary for some of the arguments in the proof of Theorem 3.2. We can show that even with larger budgets, if we restrict attention to profiles $p\in[0,1)^{n}$, an equilibrium is maximal if and only if it is equilibrium of an SGE contract.
To illustrate why we need to restrict attention to interior equilibria in the larger budget case, we provide here an example where an SGE contract induces a non-maximal equilibrium, and an example where a maximal equilibrium is induced by a non-SGE contract. First, consider the SGE contract which pays the entire budget to the successful agent with the lowest index. If $B>c_{1}^{\prime}(1)$, then the unique equilibrium is for agent $1$ to choose $p_{1}=1$ and all other agents to choose $p_{i}=0$. But this is dominated by the unique equilibrium of the piece-rate contract which pays $c_{1}^{\prime}(1)$ to agent $1$ if they succeed and $B-c_{1}^{\prime}(1)$ to agent $2$ if they succeed. Thus, in general, SGE contracts may lead to non-maximal equilibria, but in such a case, the equilibrium would necessarily involve some agent choosing $p_{i}=1$. Now suppose $B>\sum_{i\in[n]}c_{i}^{\prime}(1)$. In this case, the profile where all agents choose $p_{i}=1$, which in this case is the unique maximal equilibrium, is implemented by the piece-rate contract which pays $c_{i}^{\prime}(1)$ to agent $i$ if they succeed and 00 if they fail. Thus, in general, maximal equilibria may be implemented by contracts that are not SGE, but again, any such equilibrium necessarily involves some agent choosing $p_{i}=1$.
Implementation with Luce contracts
In this subsection, we note a useful property of Luce contracts and suggest their desirability for implementation in general multi-agent contract design environments. Our assumption that the principal has access to an exclusive-use budget departs from more standard contract design models, in which rewarding the agents is costly, and the principal trades off between the direct loss from rewarding the agents and the gain from inducing the agents to exert more effort. In such environments, where the principal is not budget-constrained, the principal has access to a larger set of contracts for the purpose of implementing any profile $q\in[0,1]^{n}$.
[…]
If $q$ satisfies (5), so that there is a Luce contract $f$ that implements $q$ (Proposition 4.2), it turns out that the total payment under any other FGN contract $g$ that implements $q$ is a mean-preserving spread of the total payment under $f$ [^2] In the special case where all agents are homogeneous and $q$ is such that $q_{1}=q_{2}=\dots=q_{n}$, we can show that $q$ satisfies (5), and an equal-split-amongst-successful contract with some budget $B$ implements $q$. To see why, notice that under the Luce contract $f$, the total payment is 00 with probability $\Pi_{i\in[n]}(1-q_{i})$, and it is some $B>0$ with the remaining probability. Since $g$ is FGN, it must also have a total payment of 00 with probability $\Pi_{i\in[n]}(1-q_{i})$, and since the expected total payment from $f$ and $g$ is the same, with the remaining probability, the total payment must be spread around $B$ in a way that the mean total payment is preserved. It follows that Luce contracts not only minimize the maximum payment the principal would have to make, but they also minimize the variance of the total payment among all FGN contracts that implement $q$. Thus, while Luce contracts are optimal in environments where the principal is budget-constrained (Theorem 5), we believe they also provide a useful alternative mechanism for inducing effort over well-studied contracts, like piece-rate and bonus-pool contracts, even in standard environments where the principal may not be budget-constrained.
@@ 6 Conclusion
We study a multi-agent contract design problem with a budget-constrained principal. In a model with independent and binary outcomes for each agent, we introduce a novel class of contracts, called Luce contracts, and demonstrate their sufficiency for implementing the maximal set of implementable equilibria. It follows that for any objective of the principal, there is always a Luce contract that is optimal. Additionally, we highlight the relevance of Luce contracts in more standard contract design environments where the principal is not budget-constrained, as they provide a means for inducing effort while minimizing the maximum total payment and also the variance of the total payment.

### v3

@@ 4 Luce contracts without budget constraints
By Proposition 2, for any FGN contract $h$ (with budget $B$) that implements $p$, $\mathbb{E}_{p}[T_{h,B}]=\sum_{i}p_{i}\cdot c_{i}^{\prime}(p_{i})$. Moreover, since $h$ is FGN, $S=\emptyset\Rightarrow T_{h,B}=0$ and $\mathbb{E}_{p}[T_{h,B}]=\mathbb{P}_{p}[S\neq\emptyset]\cdot\mathbb{E}_{p}[T_{h,B}\,|\,S\neq\emptyset]$, so $\mathbb{E}_{p}[T_{f,B}\,|\,S\neq\emptyset]=\mathbb{E}_{p}[T_{g,B^{\prime}}\,|\,S\neq\emptyset]$.
Observe that $S\neq\emptyset\Rightarrow T_{f,B}=B$, and it follows that $\mathbb{E}_{p}[T_{g,B^{\prime}}\,|\,S\neq\emptyset]=B$. Now, let $X=0$ if $S=\emptyset$ and $X=T_{g,B^{\prime}}-B$ otherwise. Then $T_{g,B^{\prime}}=T_{f,B}+X$, $\mathbb{E}_{p}[X\,|\,T_{f,B}=0]=\mathbb{E}_{p}[X\,|\,S=\emptyset]=0$, and
$\mathbb{E}_{p}[X\,|\,T_{f,B}=B]=\mathbb{E}_{p}[T_{g,B^{\prime}}-B\,|\,S\neq\emptyset]=0,$
so $T_{g,B^{\prime}}$ is a mean-preserving spread of $T_{f,B}$ at $p$. ∎
The variance of the total payment is minimized by Luce contracts (when they exist). Notice that Luce contracts also minimize the maximum total payment the principal might have to make. In the following example, we illustrate how the reduction in this maximum payment can be significant compared to piece-rate contracts, which are the natural choice of contract in a setting with independent agents and no budget constraint.
Example.
Consider $n$ agents with identical quadratic costs, $c_{i}(p_{i})=\frac{1}{2}Cp^{2}$, and consider the profile where all agents have the same success probability, $p_{i}=q$. This profile is implemented by the piece-rate contract which gives every agent $Cq$ if they succeed and nothing otherwise, and it is also implemented by the Luce contract which splits a budget of $B=\dfrac{nCq^{2}}{1-(1-q)^{n}}$ equally among all successful agents.
While the expected payment is $nCq^{2}$ under both contracts, the maximum possible total payment the principal might have to make is $nCq$ under the piece-rate contract and $\dfrac{nCq^{2}}{1-(1-q)^{n}}$ under the Luce contract. In particular, if the principal wants there to be one successful agent in expectation, as might be the case for crowdsourcing an innovation, it would implement $q=\frac{1}{n}$, and the total payment could be as large as $C$ under the piece-rate contract, where as the maximum possible total payment approaches 00 for large $n$ under the Luce contract.
@@ 5 Conclusion
We study a contract design problem with multiple independent agents. With binary outcomes for each agent and a budget-constrained principal, we introduce a novel class of contracts, called Luce contracts, and demonstrate their sufficiency for implementing the maximal set of implementable equilibria. It follows that for any objective of the principal, there is always a Luce contract that is optimal. Additionally, we highlight the relevance of Luce contracts in the absence of budget, as they provide a means for inducing effort while reducing the uncertainty in total payments and in particular, mitigating the risk of excessively large payouts.